Solve the differential equation 2y'''+ y''- 8y- 4y = 7 - 2e-x - cos(2x).


1
Expert's answer
2021-08-16T17:06:57-0400


Let

y=yh+yp

To find the homogeneous equitation,

An auxillary equation is

"2m^3+m^2-8m-4=0"

"2m^2(m+1)-4(2m+1)=0"

"(2m^2-4)(m+1)=0"

"(m+2)(m-2)(2m+1)=0"

"m=2,-2 ,-\frac 12"

Hence ,the homogeneous solution is

"y_h=C_1e^{\frac{-x}{2}}+C_2e^{-2x}+C_3e^{2x}"

To find the particular solution,

By trigonometric identities;

"7-2e^{-x}-cos^2(2x)=7-2e^{-x}-\frac12-\frac12cos(4x)=\frac{13}2-2e^{-x}-\frac12cos(4x)"

Now,

Let

"y_p=A+Be^{-x}+Ccos(4x)+Dsin(4x)"

By differentiation;

"y_p'=-Be^{-x}-4Csin(4x)+4Dsin(4x)"

"y_p''=Be^{-x}-16Ccos(4x)-16Dsin(4x)"

"y_p'''=-Be^{-x}+64Csin(4x)-64Dcos(4x)"

By substitution into the given equation;

"-2Be^{-x}+128Csin(4x)-128Dcos(4x)+Be^{-x}-16Ccos(4x)-16Dsin(4x)+8Be^{-x}+32Csin(4x)-32Dcos(4x)-4A-4Be^{-x}-4Ccos(4x)-4Dsin(4x)=\frac{13}2-2e^{-x}-\frac12cos(4x)"

Simplify;

"-4A+3Be^{-x}+(160C-20D)sin(4x)+(-160D-20C)cos(4x)=\frac{13}2-2e^{-x}-\frac12cos(4x)"

By comparison;

"-4A=\frac{13}2"

"A=-\frac{13}8"

"3B=-2"

"B=-\frac23"

"(160C-20D)=0"

"C=\frac18D" ....(i)

Also,

"-160D-20C=-\frac12" ...(ii)

Combining (i) and (ii) ,

"D=\frac {1}{325}"

"C=\frac{1}{2600}"

Substitute the constants into yp;

"y_p=-\frac{13}8-\frac23e^{-x}+\frac1{2600}cos(4x)+\frac1{325}sin(4x)"

Hence the general solution of the question is

"y=y_h+y_p"

"y=C_1e^{-\frac x2}+C_2e^{-2x}+C_3e^{2x}-\frac{13}2-\frac23e^{-x}+\frac{1}{2600}cos(4x)+\frac1{325}sin(4x)"